You make a one-time contribution of $6,000 today to your traditional IRA earning 8.37% per year.  You plan to make a full withdraw in 34 years. If you are in the 24% tax bracket then, how much will you be able to withdraw after taxes? Round your answer to two decimal places (Ex. $0.00).
Answer #1

Amount after 34 years = value invested * ( 1 + int rate )^ no of years = 6000 *1.0837^34 = 92,269.84

Amount after tax = 92269.84 * ( 1 - tax rate ) = 92269.84 * 0.76 = 70125.07